Resources Hint Check A of 29> What is the slope of the given demand curve? (5, $15) 15 Demand (13, $11 8 10 0 10 15 Quantity about us careers privacy policy terms of use contact us | help
Answer #1

let (x1 , y1 ) = ( 13,11) and (x2 , y2 ) = (5,15 )

so slope of demand curve ( straight Line) = (y2- y1 )/(x2- x1 )

putting the values from the above points

slope = (15- 11 ) / ( 5- 13 )

= 4/(-8 )

=  -1/2

= - 0.5 is the slope of the demand curve .

> should be negative, its a decrasing slope

declan mckeown Fri, Jan 14, 2022 12:44 PM